A 10 ft plank is leaning against a wall. If at a certain instant the bottom of the plank is 2 ft from the wall and is being pushed toward the wall at the rate of 6 in/s , how fast is the acute angle that the plank makes with the ground increasing?


1
Expert's answer
2021-10-27T11:01:09-0400

Solution:

"\dfrac{dx}{dt}=-6\ in/sec=-0.5\ ft/sec\ ...(i)
\\ \dfrac{d \theta}{dt}=? 
\\ \cos \theta=\dfrac{x}{10}"

On differentiating both sides w.r.t "t",

"-\sin \theta \dfrac{d \theta}{dt} =\dfrac1{10}\dfrac{dx}{dt}" ...(ii)

Using pythagoras theorem,

"h^2 +x^2 =10^2"

When "x=2"

"h^2 +2^2 =10^2
\\ \Rightarrow h^2 =100-4=96
\\ \Rightarrow h=\sqrt{96}=4\sqrt 6\ ft"

Now, "\sin \theta=\dfrac{h}{10}=\dfrac{4\sqrt6}{10}=\dfrac{2\sqrt6}{5}" ...(iii)

Using (i) and (iii) in (ii),

"-(\dfrac{2\sqrt6}{5}) \dfrac{d \theta}{dt} =\dfrac1{10}\times (-0.5)
\\ \Rightarrow \dfrac{d \theta}{dt}=0.051\ rad/sec"